OpenStudy (anonymous):

The width of a retangle is fixed at 21cm. What length will make the perimeter greater then 76cm?

OpenStudy (anonymous):

ok well I have go this far I think. a retangle has four sides. two sides are fixed at 21cm so that would be 42cm. they want it to be greater then 76.

OpenStudy (anonymous):

if you substract 76-42=34 divide by 2 =17 so can 17 be the answer to the question?

OpenStudy (anonymous):

or it could be 18 cause that is larger also. so it would be 42+36= 78

OpenStudy (anonymous):

Go slow. the width is 21, therefore 2(21)

OpenStudy (anonymous):

Call the length x, therefore 2x

OpenStudy (anonymous):

2*21=441

OpenStudy (anonymous):

2w + 2l > 76 This is the general equation, use your numbers from what I posted above to put in this equation.

OpenStudy (anonymous):

441>76

OpenStudy (anonymous):

the width is 21*2=441

OpenStudy (anonymous):

so would that be 441+?=

OpenStudy (anonymous):

You are locked in on something, and apparently ignored everything I have written.

OpenStudy (anonymous):

well you have said 2w which is = width correct

OpenStudy (anonymous):

The width is 2(21) that is given but that is not equal to 441?! That is mathematical nonsense.

OpenStudy (anonymous):

I am trying to understand. Ok well I took it as 2*21 sry!!!!!

OpenStudy (anonymous):

The perimeter is twice width and twice length. So the equation is 2(21)+2x>76

OpenStudy (anonymous):

that is 42 sry I wrote it wrong.

OpenStudy (anonymous):

That is the equation, in the equation x represents the length. Depending on what answer your teacher wants you present the equation or you solve for x if you specifically want the length.

OpenStudy (anonymous):

it just say the lenght must be greater then ? cm.

OpenStudy (anonymous):

that is why I am came up with the answer that I have had.

OpenStudy (anonymous):

I gotten 2x+42 so far

OpenStudy (anonymous):

Solve the equation and the answer x>17. You got this same answer earlier so you are apparently doing the right thing. I have just put in a neat equation so that if you have to explain it to your little brother he would understand it.

OpenStudy (anonymous):

so the answer is 17 for the lenght correct

OpenStudy (anonymous):

You don't have to ask me, just solve the equation for x; and remove all doubt. But your language and phraseology is wrong: the length must be GREATER THAN 17.

OpenStudy (anonymous):

I know sorry I am not good at this but I got the correct answer.

OpenStudy (anonymous):

Good job